Let f be a differentiable function on [alpha and beta ] and x =[alpha and beta]. Show that, if f '(x) =0 and f''(x) > 0, then f must have a local maximum at x.


1
Expert's answer
2021-05-24T02:27:02-0400

Suppose "f" has a local maximum at "x_0 ∈ (\alpha, \beta)" . For small (enough) h, "f(x_0 + h) \u2264 f(x0)."


If "h > 0" then


"\dfrac{f(x_0 + h) − f(x_0)}{ h} ≤ 0."


Similarly, if "h < 0" , then


"\dfrac{f(x_0 + h) − f(x_0)}{ h} ≥ 0."


By elementary properties of the limit, it follows that "f'(x_0) = 0."